Prove that in ∆=r^2 cotalpha/2 cot beta/2 cot gamma/2
1
Expert's answer
2019-04-19T10:34:23-0400

Consider a triangle ABC that has an incircle of radius r with center O. The incircle is tangent to AB in point P. Since the radius is perpendicular to AB, APO is right angle. Thus r is an altitude of of BOA.



Hence its area is

"\Delta_{BOA}=\frac{1}{2}rc."

By analogy, the triangles CAO and CBO have respective areas:


"\Delta_{CAO}=\frac{1}{2}rb, \space\space\space \Delta_{CBO}=\frac{1}{2}ra,"

and the total area is


"\Delta=\frac{1}{2}r(a+b+c)."

Now look a bit closer at small triangles like APO or COD or GOB. They are right-angled triangles and AP in APO for instance is:


"AP=r\text{cot}(\alpha/2)."


The area of AOP:


"\Delta_{APO}=\frac{1}{2}AP\cdot r=\frac{1}{2}\cdot r\text{cot}(\alpha/2)\cdot r=\frac{1}{2}r^2\text{cot}(\alpha/2)."

We have 6 small right-angled triangles in total and 3 of them are similar, for instance, APO and ADO. By analogy repeat the calculations for the triangles POB (=BOG) or DOC (=GOC).

The area of quadrilateral APOD is:


"A_{APOD}=r^2\text{cot}(\alpha/2)."

The total area of the triangle ABC is a sum of areas of 3 quadrilaterals (or 3 pairs of small triangles):


"\Delta=r^2 [\text{cot}(\alpha/2)+\text{cot}(\beta/2)+\text{cot}(\gamma/2)]."